OpenStudy (anonymous):

54-6/2+6=? How do I do this????

OpenStudy (anonymous):

8/8 = 8 divided by 8 which is now 1

OpenStudy (anonymous):

what? that makes no sense

OpenStudy (anonymous):

Firstly tell me what the question actually:?? \[54 - \frac{6}{2} + 6\] Right??

OpenStudy (anonymous):

Or it could be (54-6)/(2+6)

OpenStudy (anonymous):

@mattdeshurst

OpenStudy (anonymous):

Or: \[54 - \frac{6}{2+6}\]

OpenStudy (anonymous):

What is the question @mattdeshurst

OpenStudy (anonymous):